OpenStudy (anonymous):

Interesting problem: Find all positive integers such that they are equal to the square of the sum of their digits in base 10 representation. Brute force for up to \(n=25\) gives two solutions so far, \(n=1\) and \(n=9\), since \[n=1~~\Rightarrow~~1^2=1=n\\ n=9~~\Rightarrow~~9^2=81~~\Rightarrow~~8+1=9\]

OpenStudy (anonymous):

your first q and ur 99?????????????????????

OpenStudy (anonymous):

Third, actually.

OpenStudy (anonymous):

Just because he answers questions doesn't mean he can't ask them.

OpenStudy (anonymous):

im not sayin that ... like really

OpenStudy (anonymous):

I don't think your math matches your question description above.

OpenStudy (anonymous):

Suppose \(f(x)\) is the sum of the digits of \(x\) written in base \(10\). You found \(f(n^2)=n\), but the question seems to ask \(n=[f(n)]^2\).

OpenStudy (anonymous):

One thing that could help is \[ 1\leq f(n)\leq 9(\lfloor\log_{10}(n)\rfloor+1) \]

OpenStudy (anonymous):

It will help you figure out if there are any boundaries for solutions.

OpenStudy (anonymous):

In this case \(\lfloor \log_{10}(n)\rfloor + 1\) is the number of digits that \(n\) has. The highest it can be is for all of them to be \(9\).

OpenStudy (anonymous):

Oh I see, I misinterpreted the question... Thanks for clearing that up @wio